# Weak-Jamming Detection in IEEE 802.11 Networks: Techniques, Scenarios and Mobility

### v1

@@ 9 Conclusion and Future Work
In this paper, we have presented an experimental evaluation of weak-jamming detection solutions, i.e., applying when the power of the jammer at the receiver is low, and the Bit-Error-Rate of the communication link is still not affected by the jammer. To this aim, we have extended our previous solutions to make them work reliably with an actual communication technology used by Commercial Off-The-Shelf devices, i.e., IEEE 802.11, and with the complex modulation schemes used in such a standard, i.e., OFDM with BPSK, QPSK, 16-QAM, and 64-QAM. We have also enhanced the state-of-the-art methodology for jamming detection through a new image generation approach, generalizing better to complex real-world modulation schemes. Thanks to an extensive experimental campaign using low-cost SDRs, we demonstrated the effectiveness of our enhanced solution in three heterogeneous wireless environments (indoor, outdoor with multipath, outdoor with minimal multipath) across an extensive range of jamming and communication parameters. As part of our future work, we plan to deploy our solution on real COTSs to evaluate the computational burden and energy overhead it incurs on such constrained devices.

### v2

@@ 9 Conclusion and Future Work
In this paper, we have presented an experimental evaluation of weak-jamming detection solutions, i.e., applying when the power of the jammer at the receiver is low, and the Bit-Error-Rate of the communication link is still not affected by the jammer. To this aim, we have extended our previous solutions to make them work reliably with an actual communication technology used by Commercial Off-The-Shelf devices, i.e., IEEE 802.11, and with the complex modulation schemes used in such a standard, i.e., OFDM with BPSK, QPSK, 16-QAM, and 64-QAM. We have also enhanced the state-of-the-art methodology for jamming detection through a new image generation approach, generalizing better to complex real-world modulation schemes. Thanks to an extensive experimental campaign using low-cost SDRs, we demonstrated the effectiveness of our enhanced solution in three heterogeneous wireless environments (indoor, outdoor with multipath, outdoor with minimal multipath) across an extensive range of jamming and communication parameters. As part of our future work, we plan to deploy our solution on real COTSs to evaluate the computational burden and energy overhead it incurs on such constrained devices. We will also investigate to what extent a sudden change in the movement type affect jamming detection capabilities.